# ToMCCA-3: A realistic 3-body coalescence model

### v1

@@ 3 ToMCCA upgrade for $A=3$ coalescence
For the extension of the coalescence predictions, the ToMCCA model [36] had to be upgraded to incorporate the production of $A=3$ nuclei. The first step is to include an additional particle loop. For 3He this loop adds another proton, looping over $N_{\mathrm{p}}-1$ nucleons. Due to the isospin symmetry assumed in ToMCCA, the case of 3H is also covered since the first and third particles can be assumed to be neutrons. The case of ${}^{3}_{\Lambda}$ is slightly more complicated. Here, the third loop includes a $\Lambda$, which no longer abides by the isospin symmetry. The yields are modified by the $\Lambda$/p ratio measured by the ALICE collaboration as a function of multiplicity [54]. Due to the very similar mass of $\Lambda$ and proton, modifications of the $p_{\mathrm{T}}$ spectral shape can ne neglected. The second major change is the used phase space of nucleons. Previously, ToMCCA provided an uncorrelated phase space, including only very basic correlations between particles, which were mainly driven by angular correlations. However, fully-fledged Monte Carlo generators include more detailed correlations, especially between the spatial position and momentum of each particle. In order to incorporate these into ToMCCA, the EPOS 3444The EPOS 3.117 event generator was used with the special setting to force the impact parameter $b=0$. event generator was utilized, and the two-particle phase space was extracted. This means that for each nucleon-nucleon pair in the EPOS simulation, the relative momentum $k^{*}$, distance, and $\langle m_{\mathrm{T}}\rangle$ were extracted and stored. Further, each event was classified by the mid-rapidity multiplicity $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$. This provides a 4-dimensional distribution encoding the entire two-particle correlations provided by EPOS. In ToMCCA, each pair draws its source size from this distribution once $k^{*}$, $N_{\mathrm{ch}}$ and $\langle m_{\mathrm{T}}\rangle$ are determined. The phase space is assumed to be equal for nucleon-nucleon and nucleon-hyperon pairs. This leaves three different source sizes for the nucleon triplet. These three sizes are averaged according to
@@
which has been obtained by comparing the mean of the Gaussian source distribution of three equal-sized sources to one of three independent sources. This change of the phase space also requires a refitting of the source size. The source size scaling with $\langle m_{\mathrm{T}}\rangle$ has been previously parameterized as a power law
$\sigma(\langle m_{\mathrm{T}}\rangle)=B(\langle m_{\mathrm{T}}\rangle)^{C}.$ (42)
The parameters $B$ and $C$ also depend on $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ 555for sake of readability $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $\rightarrow N_{\mathrm{ch}}$
$B(N_{\mathrm{ch}})$ $=\frac{0.693N_{\mathrm{ch}}^{1/3}}{1+e^{-0.2134(N_{\mathrm{ch}}-15.4682)}}-\frac{1.6483}{1+e^{-0.2134(N_{\mathrm{ch}}-15.4682)}}+0.9234,$ (43)
$C(N_{\mathrm{ch}})$ $=\frac{0.6152}{0.456+e^{-2.9428N_{\mathrm{ch}}+2.7995}}+0.1525.$ (44)
@@
where $\Delta\varphi_{12}=\varphi_{1}-\varphi_{2}$ and $a=1,b=\pi/2$. The parameters $c$ and $N$ depend on the mean charged particle multiplicity of the event $N_{\mathrm{ch}}$
$c=$ $0.8095+0.1285\cdot N_{\mathrm{ch}}^{0.111}$ (46)
$N=$ $0.406\cdot N_{\mathrm{ch}}^{-0.373}.$ (47)
These parameters were obtained from fitting the $\Delta\varphi$ correlation functions as a function of multiplicity obtained by the ALICE collaboration [55]. The third particle undergoes the same procedure, but this time depending on two relative angles $\Delta\varphi_{13}$ and $\Delta\varphi_{23}$
@@ 4 Comparison to experimental data
In this section the output of ToMCCA is compared to experimental data. The AV18+UIX wave function with a momentum cutoff of $\Lambda_{c}=1$ GeV is considered as a reference, but also other wave functions are tested and differences are discussed. Figure 1 shows the 3He $p_{\mathrm{T}}$ spectra measured by ALICE [56] in four different $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ intervals: i) high multiplicity (HM, blue) with $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $=31.5\pm 0.3$, ii) minimum bias (MB, orange) with $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $=6.9\pm 0.1$, and a subdivision of minimum bias into iii) MB-I (red) and iv) MB-II (purple), with $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $=18.7\pm 0.3$ and $6.0\pm 0.2$ respectively. The shown experimental uncertainties are the statistical and systematic uncertainty added in quadrature. For the minimum bias predictions, the published multiplicity distributions [57] were used instead of the Erlang functions used for the multiplicity-dependent studies. The predictions from ToMCCA using the AV18+UIX wave function are shown as colored lines. Overall, an excellent agreement within $2\sigma$ is reached with all data points.
@@
Figure 1: (left) The 3He $p_{\mathrm{T}}$ spectra measured by ALICE [56] in pp collisions at $\sqrt{s}=13$ TeV in four intervals of multiplicity. The corresponding predictions for ToMCCA are shown as the colored bands. (right) The 3He/p ratio as a function of $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ and comparison to the ALICE measurements [56]. Different assumptions for the wave function are tested, only 2-body forces based on the Argonne $\text{Argonne }v_{18}$ potential (blue), 2-body+3-body Argonne $\text{Argonne }v_{18}$ +UIX wave function for 3He and 3H as well as wave function based on the 2-body Minnesota potential. The predictions for 13 TeV minimum bias collisions is shown as the hatched box.
@@
Reference [58] predicted that there should be a difference in production yield between 3He and 3H due to their different size ($r_{\mathrm{{}^{3}He}}/r_{\mathrm{{}^{3}H}}\approx 1.11$), which gets more pronounced towards lower $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ corresponding to smaller source radii. The interplay between the size of the nucleus ($r^{3}\text{He}=1.77$ fm, $r^{3}\text{H}=1.6$fm) and the source size ($\sigma\approx 1-1.5$ fm) causes a suppression of larger nuclei for smaller source sizes. The ToMCCA prediction of the yield ratio between 3H and 3He is shown in Fig. 2 (left) alongside the previously mentioned prediction and the only measurement in small systems by ALICE [56] in pp collisions at $\sqrt{s}=13$ TeV. Interestingly, ToMCCA predicts no dependence of this ratio on the source size, showing instead a constant ratio of $0.995\pm 0.007$ over the whole multiplicity range. The main difference between the two predictions are the nuclear wave functions. While Ref. [58] uses Gaussian wave functions, ToMCCA uses more realistic assumptions on the structure of the nuclei.
Figure 2: (left) 3He/3H ratio as a function of multiplicity measured by ALICE in high-multiplicity pp collisions at $\sqrt{s}=13$ TeV. The purple band shows the predictions by ToMCCA using 2+3-body interaction in both cases, the 3He and the 3H. The dashed and dot-dashed lines indicate predictions by  [58] using different assumptions of a Gaussian wave function. (right) The ratio of 3He using only the AV18 wave function over 3He using the AV18+UIX wave function as a function of multiplicity. Further the ratios of 3He and 3H obtained using the Minnesota potential to the AV18+UIX case are shown.
@@
Figure 2 (right) shows the effect of the 3-body force on the 3He and 3H yields. The ratio of the 3He yields between AV18, which considers only the 2-body forces, and AV18/UIX, which considers 2 and 3-body forces (purple band), is close to unity for very small $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$, with an increasing deviation towards higher multiplicities, up to 18% at $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}=45$. This shows an enhancement of the yield when including attractive 3-body forces, reducing the size of the nucleus. The effect even gets more pronounced at larger multiplicities, corresponding to larger source sizes. This goes against the findings of Ref. [58], where the yield of the smaller nucleus (e.g. 3H) compared to the larger nucleus (e.g. 3He) is enhanced at low multiplicities and smaller source sizes. Similarly to the case of 3H/3He, the size ultimately seems to not be the driving factor in nuclear production but the underlying interactions between the nucleons. Indeed, the UIX potential is mainly driven by long-range three-nucleon two-pion exchange. At larger sources, the long-range effects of the interaction become more important, and the attractive nature of the 3-body force increases the yields. Studies with other 3-body forces with varying ranges could be performed to quantify the effect of the interplay between source size and 3-body interaction range. This is relegated to future studies. Another realistic assumption for the nuclear wave function is given by the Minnesota potential [40], a simplified version of the nucleon-nucleon interaction (it does not include any tensor component). Despite its simplicity and the absence of a 3-body component, it is able to reproduce the binding energy of 3He and 3H within 1%. The ratio to the AV18+UIX wave function predictions shows that this potential is also equally able to reproduce the yields, with a ratio of $0.984\pm 0.016$ in the case of 3He and $0.960\pm 0.027$ in the case of 3H. Interestingly, the 3H yield is systematically lower by $\sim(2.5\pm 2)\%$, which is, considering the model uncertainties (Sec. 4.3), not significant. A similar result can be seen in Fig. 1(right). It shows the 3He/p ratio as a function of $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ compared to the measurements by ALICE [56]. The ToMCCA predictions for all wave functions are shown as the colored lines and the result for the true minimum bias prediction using AV18+UIX is shown as the hatched box. All wave functions except the AV18 only assumption reproduce the measurements.
4.2 ${}^{3}_{\Lambda}$ production
@@
ToMCCA is extended to include hyperon production as described in Sec. 3. For the ${}^{3}_{\Lambda}$, the Congleton wave function is used (see Sec. 2.2). Figure 3 (left) shows the ${}^{3}_{\Lambda}$ $p_{\mathrm{T}}$ spectra for pp collisions in the same multiplicity classes as Fig. 1. No published spectra are available for comparisons at this time for pp collisions. Figure 3 (right) shows the ${}^{3}_{\Lambda}$/3He ratio as a function of $p_{\mathrm{T}}$. This observable is sensitive to differences between thermal production models and coalescence models. Thermal models usually determine the $p_{\mathrm{T}}$ spectra using blast-wave parameterizations, mimicking radial flow effects. Radial flow causes heavier particles to have harder spectra, which causes this ratio to rise with $p_{\mathrm{T}}$. On the other end, in coalescence models, the driving factor is the source size. At larger $p_{\mathrm{T}}$ a smaller source size is observed. This causes suppression of larger nuclei such as ${}^{3}_{\Lambda}$, and hence the ratio is suppressed. It is unclear how this observation stands compared to the 3H/3He ratio, where the size of the nuclei did not influence the result.
@@
Figure 3: (left) The ${}^{3}_{\Lambda}$ $p_{\mathrm{T}}$ spectra using the Congleton wave function predicted by ToMCCA for the same multiplicities in which the 3He measurements were performed in pp collisions at $\sqrt{s}=13$ TeV. No experimental data is available for comparison at this time. (right) Ratio of ${}^{3}_{\Lambda}$/3He as a function of $p_{\mathrm{T}}$ for various multiplicity classes as predicted by ToMCCA using the Congleton wave function for ${}^{3}_{\Lambda}$ and Argonne $\text{Argonne }v_{18}$ +UIX for 3He.
Figure 4 (left) depicts the ratio of ${}^{3}_{\Lambda}$/$\Lambda$ as a function of $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$. Similarly to the $p_{\mathrm{T}}$ spectra, no integrated yields have been published for pp collisions yet. Only one measurement from p-Pb collisions is shown, however it is unclear if p-Pb and pp results should be compatible. The ToMCCA prediction is shown as the purple band. The purple box with the hatching shows the result of using the minimum bias multiplicity distribution, as used for 3He. The yield is $\approx 3$ times larger than using the Erlang distribution for $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $=6.9$, due to the much larger width of the distribution. Since coalescence is a non-linear process, i.e., it does not depend on the number of nucleons but the number of nucleon pairs, it is sensitive to higher moments of the multiplicity distribution. The right side of Fig. 4 shows the strangeness population factor $S_{3}=\frac{{}^{3}_{\Lambda}\mathrm{H}\times\mathrm{p}}{{}^{3}\mathrm{He}\times\Lambda}$. For the 3He predictions, the AV18+UIX was used. The predictions from Ref. [58] are shown as the dashed and dot-dashed lines. For $S_{3}$ the 3He yield was obtained using the 3-body coalescence approach in both cases.
Figure 4: (left) ${}^{3}_{\Lambda}$/$\Lambda$ ratio predicted by ToMCCA using the Congleton wave function are shown as the purple band. The dashed and dot-dashed lines are predictions from [58] using different assumptions for a Gaussian wave function. (right) The purple band shows the predictions of the $S_{3}$ parameter by ToMCCA using the Congleton wave function. The dashed and dot-dashed lines are the same model as on the left [58].

### v2

@@ 3 ToMCCA upgrade for $A=3$ coalescence
For the extension of the coalescence predictions, the ToMCCA model [42] had to be upgraded to incorporate the production of $A=3$ nuclei. The first step is to include an additional particle loop. For 3He this loop adds another proton, looping over $N_{\mathrm{p}}-1$ nucleons. Due to the isospin symmetry assumed in ToMCCA, the case of 3H is also covered since the first and third particles can be assumed to be neutrons. The case of $^{3}_{\Lambda}$H is slightly more complicated. Here, the third loop includes a $\Lambda$, which no longer abides by the isospin symmetry. The measured $\Lambda$ spectra from the ALICE collaboration [2] are parameterized similarly to the protons and included. The spectra are characterized using a Levy-Tsallis function
$\frac{\mathrm{d}^{2}N}{\mathrm{d}y\mathrm{d}p_{\mathrm{T}}}=\frac{p_{\mathrm{T}}\frac{\mathrm{d}N}{\mathrm{d}y}(n-1)(n-2)}{nT(nT+m_{\Lambda}(n-2))}\left(1+\frac{\sqrt{m_{\Lambda}^{2}+p_{\mathrm{T}}^{2}}-m_{\Lambda}}{nT}\right)^{-n}\ ,$ (37)
where $m_{\Lambda}=1.115\mathrm{\penalty 10000\ GeV}/c^{2}$ is used and the parameters are described as a function of the mid-rapidity multiplicity $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$  [^5] For sake of readability $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $\rightarrow N_{\mathrm{ch}}$ as follows
$\frac{\mathrm{d}N}{\mathrm{d}y}$ $=\frac{N_{\mathrm{ch}}^{2}}{0.7597+N_{\mathrm{ch}}^{2}}0.03646N_{\mathrm{ch}}$ (38)
$T$ $=0.08306N_{\mathrm{ch}}^{0.5094}+0.04290$ (39)
$n$ $=0.3050N_{\mathrm{ch}}+6.331,$ (40)
and the resulting $\Lambda$-spectra from ToMCCA are compared to the measurement in Fig. 6.
The second major change is the used phase space of nucleons. Previously, ToMCCA provided an uncorrelated phase space, including only very basic correlations between particles, which were mainly driven by angular correlations. However, fully-fledged Monte Carlo generators include more detailed correlations, especially between the spatial position and momentum of each particle. In order to incorporate these into ToMCCA, the EPOS 3666The EPOS 3.117 event generator was used with the special setting to force the impact parameter $b=0$. event generator was utilized, and the two-particle phase space was extracted. This means that for each nucleon-nucleon pair in the EPOS simulation, the relative momentum $k^{*}$, distance, and $\langle m_{\mathrm{T}}\rangle$ were extracted and stored. Further, each event was classified by the mid-rapidity multiplicity $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$. This provides a 4-dimensional distribution encoding the entire two-particle correlations provided by EPOS. In ToMCCA, each pair draws its source size from this distribution once $k^{*}$, $N_{\mathrm{ch}}$ and $\langle m_{\mathrm{T}}\rangle$ are determined. The phase space is assumed to be equal for nucleon-nucleon and nucleon-hyperon pairs. This leaves three different source sizes for the nucleon triplet. These three sizes are averaged according to
@@
which has been obtained by comparing the mean of the Gaussian source distribution of three equal-sized sources to one of three independent sources. This change of the phase space also requires a refitting of the source size [^7] The compatibility of this new parameterization with ALICE $\sqrt{s}=5$ TeV pp data is shown in Fig. 5.. The source size scaling with $\langle m_{\mathrm{T}}\rangle$ has been previously parameterized as a power law
$\sigma(\langle m_{\mathrm{T}}\rangle)=B(\langle m_{\mathrm{T}}\rangle)^{-C}.$ (42)
The parameters $B$ and $C$ also depend on $N_{\mathrm{ch}}$ as
$B(N_{\mathrm{ch}})$ $=\frac{0.7359N_{\mathrm{ch}}^{1/3}}{1+e^{-0.1808(N_{\mathrm{ch}}-12.19)}}-\frac{1.945}{1+e^{-0.1808(N_{\mathrm{ch}}-12.19)}}+1.127,$ (43)
$C(N_{\mathrm{ch}})$ $=-\frac{0.2218}{0.3168+e^{-0.1834N_{\mathrm{ch}}-5.571}}+0.9757.$ (44)
@@
where $\Delta\varphi_{12}=\varphi_{1}-\varphi_{2}$ and $a=1,b=\pi/2$. The parameters $c$ and $N$ depend on $N_{\mathrm{ch}}$ like
$c=$ $0.8095+0.1285\cdot N_{\mathrm{ch}}^{0.1114}$ (46)
$N=$ $0.4061\cdot N_{\mathrm{ch}}^{-0.3725}.$ (47)
These parameters were obtained from fitting the $\Delta\varphi$ correlation functions as a function of multiplicity obtained by the ALICE collaboration [19]. The third particle undergoes the same procedure, but this time depending on two relative angles $\Delta\varphi_{13}$ and $\Delta\varphi_{23}$, giving
@@ 4 Comparison to experimental data
In this section the output of ToMCCA is compared to experimental data. The AV18+UIX wave function is considered as a reference, but also other wave functions are tested and differences are discussed. These wave functions are calculated using the PHH method [38]. Figure 1 (left) shows the 3He $p_{\mathrm{T}}$ spectra measured by ALICE [9] in four different $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ intervals: i) high multiplicity (HM, blue) with $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $31.5\pm 0.3$, ii) minimum bias [^8] INEL¿0 by ALICE definitions. (MB, orange) with $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $6.9\pm 0.1$, and a subdivision of minimum bias into iii) MB-I (red) and iv) MB-II (purple), with $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $18.7\pm 0.3$ and $6.0\pm 0.2$ respectively. The shown experimental uncertainties are statistical (errorbars) and systematic (boxes). For the minimum bias predictions, the published multiplicity distributions [13] were used (See Sec. D) instead of the Erlang functions used for the multiplicity-dependent studies [42]. The predictions from ToMCCA using the AV18+UIX wave function are shown as colored lines. Figure 1 (right) shows the integrated yield ratios 3He/p and 3H/p as a function of $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$. Shown are the measurements by the ALICE collaboration [9, 8] in pp collisions at $\sqrt{s}=5$ TeV and $\sqrt{s}=13$ TeV. Shown are the systematic uncertainties as boxes and statistical ones as errorbars. The colored lines indicate the predictions from ToMCCA using different assumptions of the wave function. The width of the line indicates the statistical uncertainty of the model. In the inlet the minimum bias (INEL>0) measurement is compared to the model predictions. Importantly, these are not to be compared to the multiplicity continuum lines, since minimum bias collisions have a very wide multiplicity distribution. While all predictions with the correct binding energy (AV18+UIX or Minnesota) reproduce the data at 13 TeV well, there seems to be a tension in the higher-multiplicity measurement at $\sqrt{s}=5$ TeV. Overall, the model reproduces all measurements within 2 standard deviations, when considering the model uncertainties (see Sec. 4.3). Measurements in intermediate mass collision systems, such as O–O and Ne–Ne, could help rule out the two-body only hypothesis for the wave function.
@@
Figure 1: (left) The 3He $p_{\mathrm{T}}$ spectra measured by ALICE [9] in pp collisions at $\sqrt{s}=13$ TeV in four intervals of multiplicity. The corresponding predictions for ToMCCA are shown as the colored bands. The width of the central line indicates the statistical uncertainty, while the shaded band shows the systematic uncertainties of $\pm 17\%$ (see Sec. 4.3). (right) The 3He/p (3H/p) ratio as a function of $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ and comparison to the ALICE measurements [9, 8]. Different assumptions for the wave function are tested, only 2-body forces based on the $\text{Argonne }v_{18}$ potential, 2-body+3-body $\text{Argonne }v_{18}$ +UIX wave function for 3He and 3H, as well as wave function based on the 2-body Minnesota potential. The predictions and the ALICE measurement for 13 TeV minimum bias collisions are shown as the squares and circles, magnified in the inlet. For the sake of visibility the points are shifted on the x-axis.
@@
Reference [48] predicted that there should be a difference in production yield between 3He and 3H due to their different size ($r_{\mathrm{{}^{3}He}}/r_{\mathrm{{}^{3}H}}\approx 1.11$), which gets more pronounced towards lower $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ corresponding to smaller source radii. The interplay between the size of the nucleus ($r_{{}^{3}\text{He}}=1.77$ fm, $r_{{}^{3}\text{H}}=1.6$fm) and the source size ($\sigma\approx 1-1.5$ fm) causes a suppression of larger nuclei for smaller source sizes. The ToMCCA prediction of the yield ratio between 3H and 3He as a function of $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ (right) and as a function of $p_{\mathrm{T}}$ (left) is shown in Fig. 2. Interestingly, ToMCCA predicts no dependence of this ratio on the source size, showing instead a constant ratio of $0.996\pm 0.004$ over the whole multiplicity range, with a small deviation from unity and a ratio of $0.991\pm 0.002$ from $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ $\in[0,10]$. The main difference between the two predictions are the nuclear wave functions. While Ref. [48] uses Gaussian wave functions, ToMCCA uses more realistic assumptions on the structure of the nuclei. Furthermore, previous work has focused on reproducing the charge (or proton) radii of nuclei rather than the mass radii. Since 3He is doubly charged, its charge radius naturally increases. From AV18+UIX calculations, one can obtain the mass radii for both nuclei, and they are equal within 2% ($r_{M,^{3}\mathrm{H}}=1.683$ fm, $r_{M,^{3}\mathrm{He}}=1.715$ fm, see Tab. 1 for details.)
Figure 2: (left) 3H/3He ratio as a function of $p_{\mathrm{T}}$ predicted by ToMCCA for high multiplicity (HM, $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =31.5) and minimum bias (MB, $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =6.9) collisions, using two different assumptions on the wave function (AV18+UIX and Minnesota). They are compared to the measurement by ALICE in HM pp collisions at $\sqrt{s}=13$ TeV. All predictions by ToMCCA are compatible with unity within 5% and with the measurement within two standard deviations. (right) Ratio of the ToMCCA predictions using several wave function assumptions to the $^{3}\text{He}$ one using AV18+UIX.
@@
Figure 2 (right) also shows the effect of the 3-body force on the 3He yield. The ratio of the 3He yields between AV18, which considers only the 2-body forces, and AV18+UIX, which considers 2 and 3-body forces, is close to 0.9 for very small $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$, with an increasing deviation towards higher multiplicities, up to 22% at $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}=45$. This indicates an enhancement of the yield when attractive 3-body forces are included, reducing the size of the nucleus. The effect even gets more pronounced at larger multiplicities, corresponding to larger source sizes. This goes against the findings of Ref. [48], where the yield of the smaller nucleus (e.g. 3H) compared to the larger nucleus (e.g. 3He) is enhanced at low multiplicities and smaller source sizes. Contrary to common wisdom, the size ultimately seems not to be the driving factor in nuclear production but the underlying interactions between the nucleons. Indeed, the UIX potential is mainly driven by long-range three-nucleon two-pion exchange. At larger sources, the long-range effects of the interaction become more important, and the attractive nature of the 3-body force increases the yields. Studies with other 3-body forces with varying ranges could be performed to quantify the effect of the interplay between source size and 3-body interaction range. This is relegated to future studies. Another realistic assumption for the nuclear wave function is given by the Minnesota potential [49], a simplified version of the nucleon-nucleon interaction (it does not include any tensor component). Despite its simplicity and the absence of a 3-body component, it is able to reproduce the binding energy of 3He and 3H within 2%. The ratio to the AV18+UIX wave function predictions shows that this potential is also equally able to reproduce the yields, with a ratio of $0.984\pm 0.016$ in the case of 3He and $0.960\pm 0.027$ in the case of 3H. Interestingly, the 3H yield is systematically lower by $\sim(2.7\pm 0.3)\%$. This might be explained by the worse description of the binding energy for 3H ($\approx 1.2\%$ lower) compared to 3He ($\approx 0.1\%$ lower), as shown in Tab. 1.
4.2 $^{3}_{\Lambda}$H production
@@
ToMCCA is extended to include hyperon production as described in Sec. 3. For the $^{3}_{\Lambda}$H, the Congleton wave function in combination with the $\text{Argonne }v_{18}$ wave function is used (see Sec. 2.2). Figure 3 (left) shows the $^{3}_{\Lambda}$H $p_{\mathrm{T}}$ spectra for pp collisions for various multiplicities between $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =5.6 and $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =45.9 as well as minimum bias collisions. No published spectra are available for comparisons at this time for pp collisions. Figure 3 (right) shows the $^{3}_{\Lambda}$H/3He ratio as a function of $p_{\mathrm{T}}$. This observable is sensitive to differences between thermal production models and coalescence models. Thermal models usually determine the $p_{\mathrm{T}}$ spectra using blast-wave parameterizations, mimicking radial flow effects. Radial flow causes heavier particles to have harder spectra, which causes this ratio to rise with $p_{\mathrm{T}}$. On the other hand, in coalescence models, the driving factor is the source size. At larger $p_{\mathrm{T}}$ a smaller source size is observed. This is supposed to cause suppression of larger nuclei such as $^{3}_{\Lambda}$H, and hence the ratio is suppressed. It is unclear how this observation stands compared to the 3H/3He ratio, where the size of the nuclei did not influence the result. Indeed, in ToMCCA, these ratios are almost independent of $p_{\mathrm{T}}$.
@@
Figure 3: (left) The $^{3}_{\Lambda}$H $p_{\mathrm{T}}$ spectra using the Congleton wave function predicted by ToMCCA for various multiplicities from $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =5.6 to $\langle\mathrm{d}N_{\mathrm{ch}}/\mathrm{d}\eta\rangle_{|\eta|<0.5}$ =45.9 as well as minimum bias collisions. No experimental data is available for comparison at this time. The light shaded bands indicate the model uncertainties (see Sec. 4.3), the dark shaded bands represent statistical uncertainties. (right) Ratio of $^{3}_{\Lambda}$H/3He as a function of $p_{\mathrm{T}}$ for various multiplicity classes as well as minimum bias collisions. as predicted by ToMCCA using the Congleton wave function for $^{3}_{\Lambda}$H and $\text{Argonne }v_{18}$ +UIX for 3He. The shaded bands indicate the combined statistical uncertainties. No model uncertainties are estimated for these ratios.
[…]